OpenStudy (anonymous):

ty''-(t+1)y'+y=t^2 y1=e^t y2=t+1 differential equations variable coefficient equations

OpenStudy (anonymous):

@primeralph if you get a chance and would like to help me again it would be appreciated

OpenStudy (anonymous):

do I have to let w=-x?

OpenStudy (anonymous):

so change t in the first to -wy''??

OpenStudy (primeralph):

@rperez36 I have to go now. @timo86m

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Loser66 could you help please? we went over wronskian method today lol

OpenStudy (loser66):

hey, Wronskian is affected directly to RHS, therefore, your LHS must have the coefficient of y" is 1. I mean you have to divide both side by t, then , recheck the solutions whether they still work or not. then apply Wronskian.

OpenStudy (anonymous):

yeah that's what I thought but dividing out the t still left me a lil confused cause then I still have t in the denom...

OpenStudy (anonymous):



OpenStudy (loser66):

divide both side by t, you have \[y"-\frac{t+1}{t}y'+\frac{y}{t}=t\] check solution: e^t , and t+1, both them work?

OpenStudy (loser66):

e^t works,

OpenStudy (anonymous):

so I just plug in the initial conditions to the standard equation? and im just supposed to choose the givens that work? he didn't give us any examples of problems like this in class :/

OpenStudy (loser66):

t+1 works, too. Now, wronskian

OpenStudy (loser66):

yes, you should make sure if you divide both side with that coefficient and solutions still work. Most of the case, they work. but when doing that, you get standard form and your g(s) changes. In your case, it turns to t, right? quite simple. do Wronskian, friend

OpenStudy (anonymous):

ok thank you I will try

OpenStudy (loser66):

surely you can do it. just that trick, \(\huge 1*y"\)

OpenStudy (anonymous):

after wronskian I got (t+1)te^2t....that was using the y1 and y2 they gave us ...or was I supposed to do it with something else?

OpenStudy (loser66):

just apply formula , nothing else hey, I need your solution from my posted to compare to mine. Can you do it for me? the whole thing a, b, c, d. Please, If you can, write it out on a paper, scan and post it. Thanks in advance

OpenStudy (loser66):

It's not in hurry, take time, 2 more days is OK

OpenStudy (anonymous):

a problem from today or yesterday? if from today I didn't see you post a problem

OpenStudy (loser66):